OpenStudy (brucebaner):

What is the missing exponent? (10^6)-------= 10^ -18

OpenStudy (brucebaner):

@DanJS

OpenStudy (danjs):

\[10^6 * 10^a = 10^{-18}\]

OpenStudy (brucebaner):

so what is A plz tell

OpenStudy (danjs):

6 + a = -18

OpenStudy (danjs):

what is a?

OpenStudy (cwrw238):

just solve that equation for a

OpenStudy (brucebaner):

okay let me solve

OpenStudy (danjs):

a = -18 - 6 \[10^6 * 10 ^{-24} = 10^{6 - 24} = 10^{-18}\]

OpenStudy (brucebaner):

is it - 24

OpenStudy (danjs):

yep

OpenStudy (brucebaner):

so the answer is 10^6*10^-24=10^-18 @DanJS

OpenStudy (danjs):

yep, remember if you are multiplying like bases raised to a power, you can add the powers together

OpenStudy (danjs):

6 - 24 = -18

OpenStudy (brucebaner):

so how would i put that in (10^6) ----- = 10^ -18

OpenStudy (danjs):

the space would be: * 10^(-24)

OpenStudy (brucebaner):

thx so much

OpenStudy (danjs):

yw